Fermi National Laboratory

Accelerator Update for July 21, 2008 - July 23, 2008


Monday July 21

The day shift began with the Tevatron (TeV) in access, with Operations monitoring the Recycler (RR) stash (292.4E10), with the Antiproton Source and the Main Injector in access, with the Booster and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) in access, with the SciBooNE / MiniBooNE (BNB) experiments taking beam, and with the NuMI off due to the MI access.

An instrumentation expert reported at 9:28 AM that he was upgrading the BPM nodes at TeV sectors A1, A2, A3, A4, and B4.

Operators prepared Booster for an access at 10:01 AM.

Operations prepared TeV F-sector for access at 10:07 AM.

RF experts reported at 10:09 AM that they were changing the modulator for MIRF18.

Power supply experts reported at 10:14 AM that they were inspecting the MI-10 lower power supply for a ground fault.

LCW experts reported at 10:36 AM that they were accessing Pbar to fix leaks.

Linac technicians reported at 11:07 AM that they were beginning maintenance on some quadrupole power supplies. This work took about two hours.

At 11:21 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 11:24 AM, KRF4 tripped off. It reset.

Operations reported at 1:48 PM that the Pbar rings, pre-target, and enclosure B accesses were complete.

Operations reported at 2:53 PM that they were having problems trying to re-establish the MI ramp, but operators re-initialized all the profiles and brought on all the power supplies (except for MI-10 lower).

Operations established beam to BNB at 3:21 PM.

FESS personnel began work on the TeV sector E1 sump pump at 3:25 PM.

Operations reported that Pbar began stacking and NuMI resumed taking beam at 3:32 PM.

Pelletron experts reported at 3:55 PM that the hardware loop for electron cooling wasn’t regulating due to a problem in the tank. After the next shot to the TeV, the experts will have to access.

At 4:23 PM, a Linac quadrupole power supply (L:QPS225) tripped off. It reset.

At 4:26 PM, an Accumulator quadrupole power supply (A:LQ) tripped off. Operations called in an expert.

At 4:35 PM, Operations reported that the Recycler had lost about 10E10 of antiprotons due to the new turn on order in the MI.

Operations prepared Pbar for a 30-minute access at 4:49 PM.

At 5:30 PM, the cryo system experts at CHL reported that replacing the TeV sector E4 flywheel didn’t go as planned; the system was contaminated. They completed the replacement by 6:17 PM and the E4 cryo system began its cool down.

Operators reported at 7:11 PM that they had searched and secured the CDF collision hall.

The Pbar power supply expert reported at 7:17 PM that A:LQ was back online.

Operations reported at 7:37 PM that everyone was out of the TeV.

At 7:46 PM, a Pbar Debuncher RF station (DRF1-3) tripped off and would not reset. Operations called in an expert.

Operations established beam to MTest at 7:59 PM.

Pbar studiers halted stacking at 8:16 PM to make measurements. Stacking resumed about 20 minutes later.

At 8:23 PM, MIRF13 tripped off for the first of three times in six minutes. It reset.

At 8:39 PM, LRF4 tripped off for the first of two times in one minute. It reset.

At 9:04 PM, LRF4 tripped off for the first of six times in 30 minutes. After the last trip, it would not reset. Operators found its anode power supply off and not responsive. Operations contacted an expert for help and at that moment the supply came back to life.

At 10:10 PM, MIRF3 tripped off. It reset.

At 10:40 PM, Operations reported that the TeV sector F2 VCB would not close. There was no indication of a problem. Operations called in an expert.

At 11:23 PM, Operations reported that they had not established beam to MT6. Operators were investigating a problem with a power supply (MT5EL).

A power supply expert reported at 11:39 PM that the F2 VCB AC controller was in lockout. He replaced a regulator for VE43.

Tuesday July 22

The midnight shift began with the TeV turn on in progress, with Operations monitoring the stash (266.7E10), with Pbar stacking (52.9E10), with BNB taking beam, and with no MTest beam due to a power supply problem.

At 12:14 AM, Operations reported that the TeV bus tripped on a QBS ramp trip due to an A3 lead failure. Operators reset the leads and began a QBS test.

At 12:41 AM, Operations reported that Pbar was not stacking well and that operators were still trying to establish beam to MTest.

Operations put the TeV through a dry squeeze at 1:02 AM.

Operations reported at 1:28 AM that operators had managed to reset MT5EL.

Operations put the TeV through a wet squeeze at 1:47 AM.

Operations put the TeV into shot setup around 2:10 AM.

Operations established store 6310 at 4 AM with an initial luminosity of 239.7E30.

At 4:35 AM, MIRF14 tripped off. It reset.

At 4:40 AM, KRF1 tripped off. It reset.

At 4:42 AM, Operations prepared the Pelletron for an access.

Operations reestablished beam to Pbar at 5:53 AM.

At 6:32 AM, MIRF14 tripped off. It reset

At 9:44 AM, Operations reported that at random times QXR wasn’t ramping properly. Operations contacted an expert who said he would investigate when MTest was turned off.

Operations disabled the MTest beam at 10 AM.

At 10:08 AM, MIRF18 tripped off and would not reset. An expert investigated and brought the station back on line 35 minutes later.

At 10:13 AM, LRF4 tripped off. A Linac expert investigated and had the station back online within 25 minutes.

Operations prepared Pbar for access at 10:35 AM.

At 10:43 AM, Operations reported that LRF4 was repeatedly tripping. Linac experts investigated and brought it back online.

A MI expert began a coalescing study at 10:48 AM.

Instrumentation experts reported at 11:39 AM that they were taking the MI flying wire system down to diagnose a problem.

At 1:09 PM, Pbar experts reported that a Pbar Debuncher beam valve (D:BV610) would not stay open. Experts were investigating.

Pbar came out of access at 3:09 PM.

Pbar resumed stacking at 3:22 PM.

A Pbar expert began a study at 3:42 PM.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons once during the day shift.

The Pbar study ended at 4:21 PM and Pbar resumed stacking.

At 4:54 PM, Operations reported that the stack was lost due to a power supply’s (A:LQ) current going to zero. (A similar problem had happened on Monday.) Operations called in a power supply expert.

The MI study ended at 5:16 PM.

A power supply expert investigating the A:LQ problem reported at 5:21 PM that its 480-volt breaker had tripped and was faulty. Operations called in other experts to look at the breaker.

MI experts reported at 6:35 PM that they found a broken bolt for the 480-volt breaker, which had caused the breaker to over-heat and trip. They completed their repairs by 7:39 PM.

Operations established beam to MTest at 8:20 PM.

At 8:49 PM, LRF4 tripped off. It reset.

D0 experimenters reported at 10:21 PM that they were beginning a one-hour luminosity study.

Wednesday July 23

The midnight shift began with Operations monitoring store 6310 (43.57E30) and the stash (134.4E10), with Pbar stacking (96.3E10), and with BNB, NuMI, and the MTest experiments taking beam.

At 3:33 AM, a Booster kicker (MP02) tripped off. It reset.

Operators reported at 4:32 AM that they had searched and secured the Pelletron.

A Pelletron expert reported at 5:31 AM that he had turned on electron cooling and was cooling the stash.

At 6:43 AM, MP02 tripped off. It reset, but operators reported some possible regulation problems.

Operations terminated store 6310 at 7:02 AM.

Operations prepared D0 for access at 7:18 AM.

At 7:33 AM, KRF6 tripped off. It reset.

The Plans for Wednesday

The plans for today are to allow D0 experimenters to complete their repairs. Operations will then turn on the TeV and allow an expert to conduct a chromaticity study during a dry squeeze. The Run Coordinator will allow Pbar experts to conduct a destructive beam study. LCW experts will clean the MI-40 heat exchanger. The Run Coordinator then wants to put in a store early this afternoon and resume normal stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/23/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory